CH13Q4

4 The following statistics are calculated by sampling from four normal populations whose variances are equal: (You may find it useful to reference the t table and the g table.) r1 = 137, n1 = 4; = 144, n2 = 4; X3 = 136, n3 = 4; 되 = 124, n4 4; MSE = 57.4 a. Use Fishers LSD method to determine which population means differ at α=0.01. (Negative values should be indicated by a minus 10 points sign. Round intermediate calculations to at least 4 decimal places. Round your answers to 2 decimal places.) eBook Hint Print References Population Mean Differences Can we conclude that the population means differ? Confidence Interval es
Answer #1

a)

critical value of t with 0.01 level and N-k=12 degree of freedom= tN-k= 3.055
Fisher's (LSD) for group i and j =                  (tN-k)*(sp*√(1/ni+1/nj)           = 16.36
Lower bound Upper bound differ
|xi-xj|-LSD |xi-xj|-LSD
-9.36 23.36 not significant difference
-15.36 17.36 not significant difference
-8.36 24.36 not significant difference
-8.36 24.36 not significant difference
3.64 36.36 significant difference
-4.36 28.36 not significant difference